# Ultrasound matrix imaging for transcranial in-vivo localization microscopy

### v1

@@ DISCUSSION
Movie S1. Result of UMI at each depth for sheep n6, acquisition 4.
@@
Acknowledgements
The authors wish to thank the laboratory Cyceron, Institut Blood and Brain (Normandie Université, INSERM) where the sheep experiments were initially performed. In particular, we would like to thank Mikael Naveau, Palma Pro Sistiaga, Romaric Saulnier, Cyrille Orset and Denis Vivien. We would also like to thank Christine Chappard for the micro-CT imaging.
Funding: This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement nos. 772786 and 610110, ResolveStroke and REMINISCENCE projects, O.C. and A.A., respectively).
Author contributions: O.C. contributed to the conception of the ULM study. A.A. initiated and supervised the UMI part of the project. A.C., F.B. and L.D. contributed to the conception of the ultrasound sequences. L.D. and A.C. contributed to data acquisition. F.B. performed the UMI analysis. L.D. and A.C. performed the ULM analysis. L.D., F.B., A.C., O.C. and A.A. contributed to the interpretation of the experimental data. F.B., L.D., A.C. and A.A. prepared the figures. F.B., L.D., and A.A. contributed to drafting the manuscript. F.B., L.D., A.C., M.F., O.C. and A.A. discussed the results and contributed to finalizing the manuscript.
Competing interests: A.A. and M.F. are inventors on a patent related to UMI held by CNRS (no. US11346819B2, published May 2022). O.C. holds patents in ultrasound super-resolution imaging. O.C. is co-founder and shareholder of the startup ResolveStroke, which holds several patents. O.C. is the co-founder of the Resolve Stroke company that commercializes a ULM platform. All authors declare that they have no other competing interests.
Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from L.D. and O.C.
Supplementary Materials
@@ Acknowledgements
The distribution of virtual sources considered for the acquisition of the reflection matrix (UMI acquisition) is displayed in Figure S1a. As described in the Methods Section, the virtual sources are located on a sphere centered on the probe midpoint. The time delays applied to each transducer to create the virtual source associated with the white spot in Fig. S1A are shown in Fig. S1B.
S2 Resolution enhancement
@@
S3 Multiple scattering rate
@@
S4 ULM images for each dataset
@@
S4 | 1 | 28/04 | 11:27 | 10:35 | 2 | 6 | S2 | ||||||||||||||||
S5 | 1 | 31/05 | 11:21 | 11:00 | 2 | 3 | S2, S4 | ||||||||||||||||
2 | 01/06 | 11:43 | 11:04 | 1.5 | 6 | 3, S2,S4 | |||||||||||||||||
S6 | 1 | 17/06 | 10:33 | 10:40 | 1.5 | 5 | S2 | ||||||||||||||||
2 | 17/06 | 10:33 | 10:58 | 2 | 5 | S2 | |||||||||||||||||
3 | 17/06 | 13:33 | 13:11 | 2.5 | 5 | S4 | |||||||||||||||||
4 | 07/07 | 11:49 | 13:00 | 1.75 | 7 | 4,6,5,7, S2, S4 | |||||||||||||||||
5 | 08/07 | 12:36 | 13:33 | 1.5 | 10 | 3,S2, S4 |

### v2

@@ DISCUSSION
Computational Insights
The computation time of the UMI process (?) is limited by the beamforming of the focused R-matrix (Eq. 1) that takes here 30 min (on GPU with CUDA language) while the aberration correction process only lasts for a few minutes. All the post-processing was realized with Matlab (R2024a) on a working station with 2 processors @2.20GHz, 128Go of RAM, and a GPU with 48 Go of dedicated memory. This computation time for UMI remains acceptable compared to the ULM process that takes 1h30min for 9000 frames. It should be noted that the UMI process (beamforming and correction) can be further speed up by reducing the number of focusing points required to extract the aberration laws.
@@
Section S2. Flow chart of the UMI process.
Section S3. Magnitude of aberration correction.
@@
Section S6 Comparison with other aberration correction methods.
Section S7 Quantitative comparison between MRA and ULM images.
@@
Fig. S2. Flow chart of the UMI process.
Fig. S3.Quantifying the magnitude of aberration correction.
@@
Fig. S6. Comparison with other aberration correction methods.
Fig. S7. Quantitative comparison between ULM and MRA cross-sections in the saggital plane.
Fig. S8. Quantitative comparison between ULM and MRA cross-sections in the coronal plane.
@@
Movie S1. Result of UMI at each depth
@@
References (?, ?).
Supplementary Material
@@
The distribution of virtual sources considered for the acquisition of the reflection matrix (UMI acquisition) is displayed in Fig. S1a. As described in the Methods Section, the virtual sources are located on a sphere centered on the probe midpoint. The time delays applied to each transducer to create the virtual source associated with the white spot in Fig. S1A are shown in Fig. S1B.
S2 Flow chart of the UMI process
S3 Magnitude of aberration correction
The impact of skull aberrations can be estimated by the Strehl ratio which directly quantifies the decrease of the PSF. Under a random phase screen model, this Strehl ratio can be shown to scale as , with , the standard deviation of the aberration phase law. However, in the present case, the skull transfer function is not of constant amplitude. It can actually be estimated through the following matrix product:
Examples of phase and modulus of the estimated transfer function are displayed in Figs. S3B and C, respectively. While the phase of is by definition equal to the aberration phase law , its modulus describes the wave amplitude across the probe aperture. As shown by Fig. S3, this amplitude is clearly not constant over the probe aperture. The evaluation of aberration correction magnitude should therefore include this amplitude variation of the aberration transfer function. To do so, a weighted Strehl ratio can be defined in order to account for both the amplitude and phase of the aberration transfer function:
This normalized Strehl ratio is an important parameter for ULM since it corresponds the correlation degree between the imaging PSF and its theoretical diffraction-limited value.
The depth-evolution of is displayed for the three sheeps in Fig. S3A. The weighted Strehl ratio is lower for sheep n6 than for sheeps n4 and 5 and 0.28, respectively). This lower Strehl ratio indicates a stronger aberration magnitude for sheep n6. This observation is in agreement with the stronger distortions exhibited by the initial RPSF for sheep n6 (Fig. 3B) and is explained by the complex structure of its skull (Fig. 3A).
S4 Resolution enhancement
@@
S5 Multiple scattering rate
@@
S6 Comparison with other aberration correction methods
The UMI framework offers the opportunity to compare our approach with other aberration correction methods. In particular, our approach shares some similarities with the iterative time-reversal (ITR) approach of speckle noise developed by Montaldo et al. (?). While matrix imaging is inspired by this work and others studies in the field of ultrasound imaging (?, ?), it features several distinct and important advances. The first is its primary building block: the broadband focused reflection matrix that precisely selects the echoes originating from a single scattering event at each depth. This time-gating operation is decisive in terms of signal-to-noise ratio since it drastically reduces the detrimental contribution of out-of-focus and multiply scattered echoes. Second, the IPR process used in the matrix imaging process is more efficient than an iterative time reversal process. Time reversal is an adaptive filter (?) that is far from being optimal for imaging since it acts as a low pass spatial frequency filter.
This detrimental effect can be highlighted by implementing the ITR approach in the UMI pipeline. To do so, the IPR process (Eq. 17) should be replaced by ITR, such that:
The time-reversal invariant corresponds to the first eigenvector of the correlation matrix and is the associated eigenvalue. The time-reversal invariant yields an estimator of the input/output transmission matrix between the probe and each focal plane:
The result of the ITR process is illustrated in Fig. S6 by applying it to the focused reflection matrix at depth mm in sheep n6, acquisition 4. Even if the phase of the time-reversal invariant (Fig. S6E) shares some similarities with the phase law resulting from the IPR process (Fig. S6B), it also exhibits random fluctuations on some parts of the probe . Those areas coincide with the low amplitude displayed the time-reversal invariant (Fig. S6D), especially on the left of the probe. This strong variation in amplitude is a manifestation of the angular filtering process induced by iterative time reversal on the received signals. When the focused reflection matrix is beamformed (Eqs. 20 and 21) using the time reversal invariant (Eq. S4), this angular filtering of ultrasound data results in enlarged focal spots, as illustrated by the resulting map of RPSFs shown in Fig. S6F. The comparison with the IPR process (Fig. S6C) therefore demonstrates the benefit of our approach compared to an ITR approach [54]. While the IPR process narrows the confocal component of each RPSF and lowers the multiple scattering background (Fig. S6C), ITR acts as an adaptive filter and enlarges each focal spot. This loss of resolution would be detrimental to ULM since it would affect the SNR and the bubble localization precision.
The same observation holds for the SVD beamformer proposed by Bendjador et al. (?). This approach is mathematically equivalent to iterative time reversal but operated from the insonification basis, i.e here the virtual source basis (?). This basis and, equivalently, the plane wave basis are actually not adequate since most of aberrations are induced by the skull heterogeneities, i.e close to the probe. Again, we can show it by implementing the SVD beamformer in the UMI framework. To do so, a distortion matrix can be built in the virtual source basis, following the same recipe as in the transducer basis (Eq. 14), and its first singular vector directly provides an estimator of the transmission matrix , such that:
Because aberrations are from being isoplanatic from the virtual source basis, the aberration phase law exhibited by is almost flat (Fig. S6H), resulting in an extremely limited aberration correction effect. Moreover, as for the ITR approach, the SVD provides an eigenvector that shows fluctuations in amplitude. It therefore acts as an angular filter on the corrected wave-field (Fig. S6G). As for ITR (Fig. S6F), the SVD beamformer results in enlarged focal spots (Figure S6I). The SVD beamformer would therefore be inefficient for improving the ULM process. Note that an even worse result would be obtained when the SVD beamformer is operated from the plane wave basis.
S7 Quantitative comparison between MRA and ULM images
For a quantitative comparison between MRA and ULM images, we will consider an inset at the junction of the posterial cerebral arteries and the Willis polygon already visible on the saggital slice displayed in Fig. 7B and D. This part of the vascular network is nicely resolved by MRA (Fig. S7D). It can therefore be used here as a ground truth for corresponding ULM images displayed in Fig. S7E and F. The effect of UMI can be quantitatively evaluated by considering the structural similarity index (SI) of each ULM image with MRA: SI=0.17 for raw ULM vs. SI=0.33 with UMI. The SI index is therefore multiplied by a factor 2 in this specific part of the image.
ULM can do better than MRA by revealing the presence of smaller vessels (?). This higher resolving power of ULM is reinforced by UMI, both in: (i) the coronal plane with the emergence of the posterior cerebral artery in Fig. S7C that was not visible in the MRA image Fig. S7A; (ii) the sagittal plane with the resolution of smaller vessels on top of the Willis polygon in Fig. S8C that are not revealed by MRA (Fig. S8A). The latter cross-section highlights the drastic gain in contrast and resolution provided by the combination of ULM and UMI (Figs. S8C and F) compared to gold standard MRA (Figs. S8A and D) and raw ULM (Figs. S8B and E). This improvement cannot be here quantified with a structural similarity index since the MRA image is of too low quality to be used as a ground truth. However, the image quality can be quantified by its entropy (?) that measures its degree of randomness. The value of is reported on each image displayed in Fig. S8. While ULM only leads to a slight decrease of by 0.2 compared to MRA (Figs. S8D,E), the entropy loss is increased by almost a factor 5 with UMI (Fig. S8F).
S8 ULM images for each dataset
@@
The image improvement can be quantified by the loss of entropy and the contrast improvement after UMI correction. Both quantities are indicated in each image of Fig. S9 for depths ranging from 20 to 50 mm. The contrast is defined as the ratio between the variance and the mean intensity of each ULM image. Those quantities are averaged over the pixels of each ULM image for which bubbles have been detected. Zero-value pixels are discarded since they would bias the estimation of entropy and contrast.
@@
| 4 | 1 | 28/04 | 11:27 | 10:35 | 2 | 6 | 3,S3,S4,S9 | ||||||||||||||||
| S5 | 1 | 31/05 | 11:21 | 11:00 | 2 | 3 | S4, S9 | ||||||||||||||||
| 2 | 01/06 | 11:43 | 11:04 | 1.5 | 6 | 3, S3, S4,S9 | |||||||||||||||||
| S6 | 1 | 17/06 | 10:33 | 10:40 | 1.5 | 5 | S4 | ||||||||||||||||
| 2 | 17/06 | 10:33 | 10:58 | 2 | 5 | S4 | |||||||||||||||||
| 3 | 17/06 | 13:33 | 13:11 | 2.5 | 5 | S9 | |||||||||||||||||
| 4 | 07/07 | 11:49 | 13:00 | 1.75 | 7 | 4,6,5,7, S3,S4, S5,S6,S9 | |||||||||||||||||
| 5 | 08/07 | 12:36 | 13:33 | 1.5 | 10 | 3,S4, S9 | |||||||||||||||||
Captions of Supplementary Movies
Movie S1: Result of UMI at each depth. Left: Selected depth. Middle Top: Local RPSFs before and after correction. Middle Bottom: Averaged RPSF before and after correction. Top right: Extracted aberration phase law. Bottom right: Resolution enhancement provided by UMI. The data shown here correspond to sheep n6 (acquisition 4)
Movie S2: ULM images for different acquisitions before and after UMI correction. 3D images are shown as maximum intensity projections.